For each of the following solutions, calculate the initial pH and the final pH after adding 0.010 mol of NaOH.

12A. For 280.0 mL of a buffer solution that is 0.205 M in HCHO2 and 0.315 M in KCHO2, calculate the initial pH and the final pH after adding 0.010 mol of NaOH.

12B. For 280.0 mL of a buffer solution that is 0.325 M in CH3CH2NH2 and 0.295 M in CH3CH2NH3Cl, calculate the initial pH and the final pH after adding 0.010 mol of NaOH.

Express your answers using two decimal places separated by a comma.
Answer #1

Note : i took Ka values from sapling . please check once whether these are match with your values or not. if any wrong that is only with Ka values. so please check before doing

12A. )

pH = pKa + log [salt / acid]

    = 3.74 + log [0.315 / 0.205]

initial pH = 3.93

when 0.010 mol NaOH is added :

pH = pKa + log [salt + C / acid - C]

    = 3.74 + log [0.315 + 0.01 / 0.205 - 0.01]

final pH = 3.96

12B.)

pH = pKa + log [base / acid]

    = 10.80 + log [0.325 / 0.295]

initial pH = 10.84

when NaOH is added :

pH = pKa + log [base + C / acid - C]

      = 10.80 + log [0.325 + 0.01 / 0.295 - 0.01]

final pH = 10.87